### v2

# DeepEdit: Knowledge Editing as Decoding with Constraints

Abstract
We propose a new perspective of knowledge editing (KE) for large language models (LLMs) that treats it as a constrained decoding problem. We design decoding constraints to regulate LLMs, ensuring coherence between reasoning steps when incorporating new knowledge. To enforce these constraints, we utilize a depth-first search to adaptively substitute new knowledge for the LLMs’ original reasoning steps, greedily seeking the optimal path of multi-hop reasoning with new knowledge. From this vantage, we propose DeepEdit: Depth-first Search-based Decoding for Knowledge Editing. DeepEdit improves the KE of LLMs by enhancing the conciseness, coherence, pertinence, and receptiveness of reasoning with new knowledge. DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions. In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which are designed to provide more precise and challenging assessments of KE approaches. Qualitatively, DeepEdit enables LLMs to produce more succinct reasoning outputs in accordance with new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
DeepEdit: Knowledge Editing as Decoding with Constraints
Yiwei Wang     Muhao Chen     Nanyun Peng     Kai-Wei Chang University of California, Los Angeles   University of California, Davis wangyw.evan@gmail.com https://wangywust.github.io/deepedit.io/
1 Introduction
Knowledge editing (KE) aims to enable large language models (LLMs) to produce outputs in accordance with the knowledge that is new or conflicted with the knowledge memorized by the LLMs. A challenging task to evaluate KE is the multi-hop question answering with new knowledge Zhong et al. (2023). In this task, LLMs should answer the questions corresponding to a chain of facts including new knowledge (see Figure 1(a)). It is essential for LLMs to deductively infer from the new knowledge in the multi-step reasoning in order to provide correct answers.
Figure 1: (a) An instance of multi-hop question answering with new knowledge. (b) The existing method MeLLo relies on the LLM generator to understand new knowledge by itself. (c) Our DeepEdit applies the constraints to the LLMs’ decoding so as to improve reasoning with new knowledge.
Figure 2: Our DeepEdit enforces LLMs to follow the decoding constraints of [Conciseness], [Coherence], [Receptiveness], and [Pertinence] so as to soundly incorporate new knowledge into LLMs’ reasoning.
Recent literature has mainly sought to perform KE through in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023a). However, in-context editing lacks direct control of the outputs; the new knowledge‘s influences are not traceable; and the performance suffers from the input-conflicting hallucinations Zhang et al. (2023b) (see Figure 1(b)).
Taking a closer look at the ground-truth output in Figure 1(a), we notice that: when reasoning with new knowledge, LLMs should orderly place new knowledge at the appropriate positions in the reasoning chain, i.e., the new knowledge should coherently bridge the adjacent reasoning steps, and avoid the memorized knowledge that is conflicted to new knowledge. This finding raises a natural question:
Can we directly control the decoding of LLMs’ to soundly incorporate the new knowledge?
The central idea of this paper is to view KE as a problem of constrained decoding. We design the decoding constraints, by following which LLMs can soundly incorporate new knowledge in LLMs reasoning. Specifically, we propose the following constraints: [Conciseness] prevents the redundant loops of reasoning; [Coherence] guarantees the coherence of adjacent steps; [Receptiveness] guarantees the LLMs’ awareness of new knowledge; [Pertinence] improves the relevance of reasoning to the input question (see Figure 4). These constraints improve LLMs’ reasoning with new knowledge from different perspectives as a whole.
Much research has proposed constrained decoding methods Meng et al. (2022c); Liu et al. (2023); Geng et al. (2023). However, classical constrained decoding methods usually require access to the token-wise distribution, which is not provided by many black-box LLMs like ChatGPT OpenAI (2022). In our work, to apply our constraints to the black-box LLMs’ decoding, we breakthrough this restriction by changing the granularity of decoding from tokens to the reasoning steps. We design all decoding actions over the step level, which bypasses the requirement of accessing the token-wise distributions.
At every iteration, we take one generated reasoning step and a few retrieved new facts as the step candidates. We verify the step candidates and filter out the ones that do not satisfy our constraints. Since there could be multiple candidates that satisfy all constraints every iteration. We operate the depth-first search that store all the valid step candidates in a stack and greedily search for the optimal path of LLMs’ multi-hop reasoning with new knowledge. We name our method as DeepEdit, the depth-first search-based decoding for knowledge editing.
The advantages of our DeepEdit on KE are three-folds. First, it does not need to retrain the LLMs like the classical model editing methods Zhu et al. (2020), which is complex and expensive. Second, the new knowledge directly contributes to the LLMs’ decoding by acting as the reasoning step candidates. This facilitates more effective reasoning with new knowledge. Third, the influence of new knowledge is monitored with decoding constraints, leading to more controllable and traceable reasoning with new knowledge. Overall, our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness.
DeepEdit can be flexibly applied to all the black-box LLMs. We evaluate DeepEdit on the KE benchmarks MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al. (2023); Meng et al. (2022a). The latter two are new benchmarks that we build to provide more precise and challenging assessments of KE methods. Qualitatively, DeepEdit leads to more succinct and faithful reasoning of LLMs that soundly incorporate new knoweldge, as shown in Figure 8. We also observe substantial quantitative improvements in the KE question-answering accuracy achieved by DeepEdit. Overall, DeepEdit improves the KE for the strong LLMs text-davinci-003, gpt-3.5-turbo-instruct, and llama2 by a significant margin OpenAI (2022); Touvron et al. (2023).
2 Related Work
Constrained Decoding
Constrained decoding Qin et al. (2020, 2022); Lu et al. (2020, 2021); Kumar et al. (2022); Liu et al. (2023); Geng et al. (2023) is a classical natural language processing topic, which is mainly applied to control the lexicon, sentiment or style of the generated text. FUDGE Yang and Klein (2021) employs weighted decoding (WD) by directly adjusting the probabilities of the vocabulary distribution with an auxiliary classifier probability. NADO Meng et al. (2022c) suggests sampling from a vocabulary distribution similar to an approximation of the sequence-level oracle. GeDi Krause et al. (2020) and DExperts Liu et al. (2021) also adopt the weighted decoding approach through the training of generative classifiers. Constrained decoding methods generally require access to the LLMs’ token-wise distribution. Different from them, our DeepEdit avoids the requirements of distribution access by decoding at the reasoning step level. We extend the applications of constrained decoding to KE. We design the semantic constraints to soundly incorporate new knowledge in LLMs’ reasoning, and verify the complex semantic constraints with LLMs.
Figure 3: Comparison of different KE methods. (upper) Model editing was the most popular KE paradigm, which edits the model parameters to update knowledge. (middle) Recently, in-context editing has become popular, which relies on the LLMs’ capability of utilizing new knowledge in the prompts. (lower) We propose DeepEdit for KE, which directly controls the LLMs’ decoding to follow new knowledge.
Knowledge Editing
Previous studies have explored multiple methods for editing the knowledge of LLMs by introducing new knowledge into static model artifacts Zhu et al. (2020); Sotoudeh and Thakur (2019); Dai et al. (2021); Hase et al. (2021); Zhou et al. (2023); Dong et al. (2022); Huang et al. (2023). Some of these approaches involve identifying and modifying model weights associated with specific concepts Meng et al. (2022a, b); Dai et al. (2021), as well as rapid adaptation facilitated by a compact auxiliary editing network Mitchell et al. (2022a); De Cao et al. (2021); Mitchell et al. (2021). Given the fast-growing parameter sizes of large language models (LLMs) Zhao et al. (2023), frequently updating LLMs with new knowledge through retraining is more and more expensive. Hence, it is vital to effectively edit the LLMs’ knowledge without retraining.
Recent work observes the superior KE performance of in-context editing. MeLLo Zhong et al. (2023) designs a single prompt to alternately conduct the text generation and new knowledge analysis using the LLM generator. Cohen et al. (2023) propose to append the new knowledge the head of the input prompt. In-context editing lacks direct control of the LLMs’ outputs and how the new knowledge influences the outputs is not traceable. Different from them, we propose the new decoding method DeepEdit for KE, which effectively augments LLMs on KE during the decoding stage. We illustrate a comparison between DeepEdit and existing KE methods in Figure 3.
Specifically, in our work, we consider a challenging setting of knowledge editing: the multi-hop question answering (QA) with knowledge editing Zhong et al. (2023). In-context learning methods combined with an optional retrieval module have become a popular approach in handling multi-hop QA tasks Press et al. (2022); Yao et al. (2022); Khattab et al. (2022); Mitchell et al. (2022b). We propose four types of decoding constraints to enhance question answering with KE. Our work improves multi-hop QA with KE by applying constraints to the reasoning of LLMs, which directly controls LLMs’ outputs to follow new knowledge.
Figure 4: Comparison between classical constrained decoding (at the token level) (upper) and our decoding at the reasoning step level (lower). Differences: 1) The token candidates is the predefined vocabulary, while the step candidates include a generated step and the retrieved facts. 2) The token candidates are ranked by generators’ outputs, while the step candidates are filtered by the constraints.
Self-refinement of Large Language Models
Some work has explored to use LLMs to revise the output of themselves in dialogue response, sentiment reversal, and mathematical reasoning Madaan et al. (2023); Shridhar et al. (2023b, a); Welleck et al. (2022); Zheng et al. (2023b); Madaan et al. (2023). However, it is difficult to rely on LLMs to accommodate the new knowledge by themselves in the self-refinement, since the new knowledge is generally conflicted with the parametric knowledge memorized by LLMs. Different from them, we do not rely on LLMs to revise the outputs by themselves. We directly control the decoding to soundly incorporate new knowledge with decoding constraints. We modify the LLMs’ outputs using knowledge retrieval and depth-first search to satisfy the constraints. We hope that our work demonstrates more potential to conduct KE for LLMs in the decoding stage.
3 Methodology
Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) Zhong et al. (2023), where LLMs should answer the questions corresponding to a chain of facts including new knowledge. This task assumes a database storing all new facts that LLMs should follow, and only a few facts can influence the ground-truth answer to a specific question.
Our KE idea is to first design the decoding constraints that can regulate LLMs to soundly incorporate new knowledge in reasoning and then enforce LLMs’ decoding to follow the constraints. In this section, we will introduce how we incorporate our idea into the design of a new KE method DeepEdit (Depth-first Search based Decoding for Knowledge Editing) to improve the KE of black-box LLMs.
3.1 Decoding at the Reasoning Step Level
Many classical constrained decoding methods Zhang et al. (2023a); Meng et al. (2022c) use an auxiliary model to adjust the token-wise distribution during decoding. They manipulate the distribution before every token’s generation to satisfy the constraints. Therefore, they require accessibility to the LLMs’ output distribution, which is not feasible for many black-box LLMs.
In this work, we change the granularity of decoding from the token level to the reasoning step level. Our decoding is designed at the step level: instead of taking the tokens predefined in a fixed vocabulary as the candidates, we take the reasoning steps as the candidates. At every iteration, we let LLMs to produce one reasoning step and retrieve a fixed number of new facts in parallel to the reasoning step as the step candidates. The new fact retrieval is based on the nearest neighbor search with the sentence-level embeddings given by BERT. The motivation of this design is that the new facts that are semantically close to the generated facts are more likely to satisfy the constraints.
Taking the single generated step and the new facts as the step candidates, we filter a valid candidate and take it as the new reasoning step. Note that there may exist more than one candidates that satisfy the constraints. We will introduce how we design the decoding constraints for KE in Section 3.2, how we verify the constraints in Section 3.3 and Section 3.4, and how we process the multiple valid step candidates to search for the optimal path of LLMs’ multi-hop reasoning in Section 3.5.
Our decoding design bypasses the requirement of accessing the token-wise distributions, since all the decoding actions are defined over the step level. Our decoding strategy is free to be applied to black-box LLMs in order to produce satisfactory outputs in accordance with new knowledge. The details are introduced next.
Figure 5: (left) At every decoding iteration, we store the reasoning step candidates (including the generated step and the retrieved knowledge) that satisfy the local constraints in a stack. (right) The global verifier determines the feasibility of the whole reasoning process by checking whether the answer is found. If not feasible, we pop the status stack to use another reasoning subprocess that satisfies the local verifier.
3.2 Designing Decoding Constraints for Knowledge Editing
To correctly answer multi-hop questions in accordance with new knowledge Zhong et al. (2023), LLMs need to generate a sequence of reasoning steps to find the answer. We propose the following constraints for the LLMs’ reasoning with KE.
  1. 1.
    [Conciseness] Every reasoning step should be different from the prior ones.
  2. 2.
    [Coherence] Every reasoning step should be relevant to the last reasoning step.
  3. 3.
    [Receptiveness] Every reasoning step cannot contradict to any new knowledge.
  4. 4.
    [Pertinence] Every reasoning step should be relevant and helpful to finding the answer.
The above constraints are not automatically satisfied by LLMs when reasoning with new knowledge. However, they are essential to helping LLMs to find the correct answers in accordance with new knowledge. Therefore, in our work, we propose a new decoding method to enforce the LLMs to follow the above constraints so as to improve the KE performance. We first analyze the constraints in detail.
Among the above four constraints, [Conciseness] guarantees that the LLMs’ reasoning is not in a loop. [Coherence] guarantees the coherence of reasoning, while [Pertinence] guarantees the relevance of every reasoning step to the question. [Receptiveness] ensures that LLMs can utilize new knowledge for reasoning. We visualize different constraints in Figure 4. Among the constraints, we regard [Conciseness], [Coherence], and [Receptiveness] as local ones: we can verify whether they are satisfied given only partial reasoning steps. The [Pertinence] constraint is regarded as a global constraint: we can verify it based on a complete reasoning process.
By following the above constraints, LLMs can orderly place new knowledge at appropriate positions in the reasoning chain and avoid the memorized knowledge that is conflicted with new knoweledge. Next, we will introduce how we verify the above constraints and control the LLMs’ decoding to satisfy the constraints.
3.3 Verifier for Local Constraints
To verify the [Conciseness] constraint, by default, whenever a new reasoning step is given, we may compare it with all the prior reasoning steps. However, this is no doubt an expensive solution at the cost of $\mathcal{O}(r^{2})$ times of verification given as many as $r$ reasoning steps.
We observe that the repeated reasoning step generally occurs in the adjacent steps. Therefore, to reduce the computational costs, we propose an approximate verification method to verify the constraint of [Conciseness]. Specifically, for every reasoning step, we only compare it with the last reasoning step, so that the whole time of doing the verification is $\mathcal{O}(r)$.
Figure 6: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k. The chosen step candidate in every iteration that satisfies all the constraints are in green. The dropped step candidates are followed by the constraints that they do not satisfy.
To satisfy the [Coherence] constraint, we need to evaluate whether the current reasoning step is relevant to the last reasoning step. However, it is challenging to directly and precisely evaluate the relevance through LLMs. We use the LLM based verifiers equipped with a hand-crafted prompt to verify whether the adjacent reasoning steps are semantically equivalent or related to verify the constraints of [Conciseness] and [Coherence] respectively.
Verifying the [Receptiveness] constraint, by default, requires determining whether the contradiction exists in the pair of the reasoning step and all the items in the basis of new knowledge. The time complexity is $\mathcal{O}(rk)$, where $k$ is the number of new knowledge items and a large value in practice. Therefore, to reduce the time complexity, we retrieve the relevant knowledge items that are semantically closest to the reasoning step and then assess the contradictions. The idea is that if two sentences are contradictory to each other, they must first be semantically relevant. In this way, we reduce the time complexity from $\mathcal{O}(rk)$ to $\mathcal{O}(r\log(k))$. Also, we design a LLM-based verifier with a hand-crafted prompt to assess whether the retrieved facts are contradictory to the generated reasoning step.
3.4 Verifier for Global Constraints
Whether a reasoning step is helpful in finding the final answer is hard to justify until the final answer is found. Therefore, we only verify the [Pertinence] constraint globally for each reasoning step given the full completed reasoning process. When the intermediate reasoning steps are not helpful but even barrier the final answer, the reasoning likely takes more steps to find the final answer. Based on this prior knowledge, we set the maximum reasoning steps as $r$, so that the reasoning is terminated when an answer is not found given more than $r$ steps. We set $r=5$ by default in our experiments.
3.5 Decoding with Depth-first Search
As we have introduced in Section 3.1, there may exist multiple step candidates that satisfy all the constraints. In this case, how can we decide which step is the next step?
To answer this question, let’s first recall how LLMs how to decide which is next token. The most simple method is the greedy decoding, which just takes the highest logit token candiate as the next token. In our case, since we do not have scores to rank step candidates but only the discrete constraint verifications to filter the candidates, we cannot simply take the “best” step as the next reasoning step. Can we have other alternate solutions?
Let’s see a more advanced token-level decoding method: beam search. Beam search takes multiple candidate tokens as the next tokens and store them in a stack to do a best-first search in the token sequence generation. We borrow the idea of the beam search to process the multiple valid step candidates with the depth-first search. Specifically, we store all the step candidates that satisfy the constraints in a stack, as shown in Figure 5. Then, in every iteration, we pop the reasoning step stored at the head of the stack as the next reasoning step and do further reasoning based on it. We name our method as DeepEdit: depth-first search based decoding for knowledge editing. We visualize our DeepEdit in Figure 1(c). Also, we present the iteration-wise outputs of our DeepEdit on an instance in Figure 6.
The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al. (2023). Different from them, our DeepEdit’s depth-first search is to enforce the LLMs’ decoding to satisfy the constraints derived for KE from a global perspective.
4 New Benchmarks for More Precise and Challenging Evaluation
We provide two new benchmarks for the evaluation of KE methods: MQuAKE-2002 and MQuAKE-hard. Both are built from the MQuAKE Zhong et al. (2023), a recent KE benchmark that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge. Every instance in MQuAKE includes a multi-hop question and one to four related edited facts, every of which can alter the ground-truth answer to the question. Zhong et al. (2023) suggests using a randomly sampled 3,000 instance subset of MQuAKE to do the evaluation, which reduces the experimental costs. This evaluation benchmark is termed as MQuAKE-3k.
4.1 Annotation Mistakes in One-Third Instances of MQuAKE-3k
There are two issues of using MQuAKE-3k to evaluate KE. The first issue is that the new knowledge from different instances MQuAKE-3k can cause conflicts and mistakes to the ground-truth answers. In other words, the ground-truth answer from instance A can be altered by the new knowledge from another instance B. We show an example of knowledge conflicts in MQuAKE-3k in Figure 7. These knowledge conflicts will make the ground-truth answers incorrect given the new knowledge that are conflicted with the answers because the inference on every instance would retrieve the new knowledge from all instances. We observe that 998 instances’ ground-truth answers are broken by the new knowledge from other instances.
4.2 New Benchmark MQuAKE-2002 for More Precise Evaluation
To address the issue of annotation mistakes in MQuAKE-3k, we provide a new subset of MQuAKE-3k, which does not have any knowledge conflict across instances. This subset includes 2,002 instances, so we term it as MQuAKE-2002. We filter out the instances of which the ground-truth answers are broken by the new knowledge from other instances to produce MQuAKE-2002. Compared with MQuAKE-3k, our MQuAKE-2002 provides a more precise evaluation for knowledge editing methods, since it remove the annotation mistakes due to knowledge conflicts across instances. The data statistics of MQuAKE-2002 is provided in Table 1.
4.3 New Benchmark MQuAKE-hard for More Challenging Evaluation
The second issue of MQuAKE-3k is that more than 66% instances in MQuAKE-3k only contain at most two edited facts that influence the answers, which are not challenging enough to evaluate the knowledge editing methods on handling multiple edited facts that can alter the ground-truth reasons. we construct a more challenging subset of MQuAKE by selecting the instances that contain the highest number of edited facts per instance. We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts. MQuAKE-hard has no overlap with MQuAKE-3k. We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts. MQuAKE-hard has no overlap with MQuAKE-3k. The data statistics of MQuAKE-hard is also provided in Table 1.
Benchmark # Instances # Hops per Instance # Edited Facts per Instance # Conflicted Instances MQuAKE-3k Zhong et al. (2023) 3,000 3.0 2.0 998 MQuAKE-2002 2,002 2.7 2.2 0 MQuAKE-hard 4,29 4.0 4.0 0
Table 1: Data statistics of different benchmarks. # Conflicted Instances represent the number of instances of which the ground-truth labels are affected by the new knowledge from other instances. An example is shown in Figure 7.
Figure 7: An example of knowledge conflicts across instances in MQuAKE-3k. New knowledge (in red) in instance 382 is conflicted with the ground-truth reasoning step (in green) of instance 22.
5 Experiments
In this section, we evaluate the KE performance of our DeepEdit method when applied to LLMs, and compare against a strong baseline MeLLo Zhong et al. (2023). Our experimental settings closely follow those of the previous work Zhong et al. (2023) to ensure a fair comparison.
5.1 Experimental Settings
We take the following KE benchmarks MQuAKE-3k Zhong et al. (2023), MQuAKE-2002, and MQuAKE-hard for evaluation. The introduction of the above benchmarks are provided in Section 4. We follow the existing work Zhong et al. (2023) to use accuracy as the main evaluation metric. For every instance, we test the accuracy of the first question in the instance, since different questions in the same instance are semantically equivalent and correspond to the same ground-truth answer. We follow Zhong et al. (2023) to set the KE batch size as 100 and full batch for evaluation. The KE batch size means the number of instances that provide new knowledge for retrieval.
Compared methods. We take the strong KE methods into comparison: MEND Mitchell et al. (2021), MEMIT Meng et al. (2022b), IKE Zheng et al. (2023a), ICE Cohen et al. (2023), and MeLLo Zhong et al. (2023). MEND produces weight updates by transforming the raw fine-tuning gradients given an edited fact. MEMIT updates feedforward networks in a range of layers to encode the edited facts. IKE and ICE conduct KE by adding new knowledge to the input prompts. MeLLo is the state-of-the-art KE method that prompts the LLMs to generate subquestions and append new knowledge to the prompts. The former two KE methods are only applicable to white-box LLMs, while the latter three can fit the block-box LLMs.
Method MQuAKE-3k MQuAKE-hard batch size of 100 full batch batch size of 100 full batch llama2-7b w/ MEND Mitchell et al. (2021) 7.2 3.9 0.5 0.2 llama2-7b w/ MEMIT Meng et al. (2022b) 7.9 4.2 0.9 0.5 llama2-7b w/ IKE Zheng et al. (2023a) 12.5 6.2 1.2 0.5 llama2-7b w/ ICE Cohen et al. (2023) 12.9 6.3 1.4 0.7 llama2-7b w/ MeLLo Zhong et al. (2023) 16.3 10.8 1.9 1.6 llama2-7b w/ DeepEdit (Ours) 27.6 11.2 14.5 7.0 gpt-3.5 w/ IKE Zheng et al. (2023a) 15.5 8.3 1.6 1.2 gpt-3.5 w/ ICE Cohen et al. (2023) 16.1 8.5 2.1 1.4 gpt-3.5 w/ MeLLo Zhong et al. (2023) 29.9 20.0 3.7 1.6 gpt-3.5 w/ DeepEdit (Ours) 47.2 38.0 51.3 48.0 text-davinci-003 w/ IKE Zheng et al. (2023a) 18.2 8.9 1.9 1.2 text-davinci-003 w/ ICE Cohen et al. (2023) 19.0 8.6 2.3 1.2 text-davinci-003 w/ MeLLo Zhong et al. (2023) 32.2 25.6 4.7 2.3 text-davinci-003 w/ DeepEdit (Ours) 49.1 38.8 52.0 49.2
Table 2: Experimental results (accuracy; %) on the dataset MQuAKE-3k and MQuAKE-hard. The batch size means the number of instances providing the edited facts for knowledge retrieval. gpt-3.5 stands for the model gpt-3.5-turbo-instruct. The best KE result on every LLM is highlighted in bold font.
Method MQuAKE-2002 gpt-3.5 w/ MeLLo 25.1 gpt-3.5 w/ DeepEdit 53.7
Table 3: Experimental results (accuracy; %) of the full-batch knowledge editing on the benchmark MQuAKE-2002.
Verifier [Conciseness] [Coherence] [Receptiveness] llama2-7b 82.3 77.9 86.3 gpt-3.5 92.8 91.5 92.2
Table 4: Experimental results (verification accuracy; %) of our verifiers for different constraints on the dataset MQuAKE-3k.
gpt-3.5 w/ DeepEdit 49.1 $\Delta$ w/o [Conciseness] 46.1 $\downarrow 6\%$ w/o [Coherence] 42.9 $\downarrow 13\%$ w/o [Receptiveness] 45.2 $\downarrow 8\%$ w/o [Pertinence] 46.6 $\downarrow 5\%$
Table 5: Experimental results (accuracy; %) on the dataset MQuAKE-3k of DeepEdit with KE batch size of 100.
Model configuration. We set the hyper-parameters of the baseline methods as suggested by their papers. We apply the KE methods to the strong LLM text-davinci-003, gpt-3.5-turbo-instruct (short as gpt-3.5), and llama2-7b (short as llama2) OpenAI (2022); Touvron et al. (2023) for evaluation. The former two are black-box LLMs and the latter one is a white-box LLM. By default, we set the decoding temperature as 0.0 to minimize the randomness of LLMs’ outputs. The DeepEdit’s verifiers use the same LLM as the generator.
Figure 8: A case study for MeLLo and our DeepEdit on the dataset MQuAKE-3k. The new knowledge are in green. We report the prediction of text-davinci-003 with MeLLo and our DeepEdit.
5.2 Overall Performance
We incorporate DeepEdit with several choices of LLMs to test its effectiveness on editing LLMs’ knowledge. We report the test accuracy on MQuAKE-3k and MQuAKE-hard in Table 2. When the KE batch size is 100, our DeepEdit improves the accuracy of question answering with KE by 69%, 58%, and 52% over MeLLo on the LLMs llama2, gpt-3.5-turbo-instruct, and text-davinci-003 respectively over the strong baseline method MeLLo on MQuAKE-3k. On MQuAKE-hard, our DeepEdit outperforms the strong baseline method MeLLo by $\times$7, $\times$12, and $\times$10 times in terms of question answering accuracy with KE on the LLMs llama2, gpt-3.5, and text-davinci-003 respectively.
The experimental results on MQuAKE-2002 is shown in Table 3. When the KE batch size is full batch, our DeepEdit still outperforms the baseline methods and improves the KE accuracy by a significant margin. Our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness. The above advantages are especially important when the LLMs have to utilize multiple new knowledge to get the correct answers, since more new knowledge requires the LLMs to be correctly aware of the new knowledge on more reasoning steps. The correct knowledge incorporation with our decoding constraints significantly improves the KE performance given more new knowledge. Overall, our DeepEdit achieves substantial improvements on KE for LLMs and consistently outperforms the baseline methods on different benchmarks.
5.3 Analysis on Verifiers
We conduct deep analysis on the different verifiers of DeepEdit. We first test the accuracy of different verifiers to verify the corresponding constraint. We randomly sample 4,389 pairs of reasoning steps from the ground-truth reasoning steps of MQuAKE as the positive pairs that satisfy the constraints and randomly pairing the same number of steps from different questions as the negative examples. We evaluate different verifiers based on llama2 and gpt-3.5 on the sampled reasoning step pairs. We report the verification accuracy in Table 4. We observe high verification accuracy of around 80% and over 90% on different constraints achieved by llama-2 and gpt-3.5 respectively. Also, we observe that the performance of verifiers based on llama2 and gpt-3.5 are positively related to the LLMs’ performance with our DeepEdit, since the effective constraint assessment forms the basis of effective reasoning of LLMs with our DeepEdit.
5.4 Ablation Study
We conduct ablation studies on DeepEdit to empirically examine the contribution of its main technical components. We evaluate the influence of different constraints on DeepEdit. We report the experimental results of the ablation study of constraints in Table 5. We observe that removing each constraint leads to substantial performance degradation in different degrees. This indicates that different constraints contribute to more effective KE from the perspectives of progress, relevance, coherence, and awareness as a whole, and removing any of them will lead to more mistakes in the reasoning of LLMs. The above results also validate that the LLMs’ outputs cannot automatically satisfy the constraints that represent the desired properties without our DeepEdit.
5.5 Case Study
Figure 8 gives a qualitative comparison example between MeLLo and our DeepEdit on MQuaKE-3k. The results show that the strong KE method MeLLo makes mistakes on KE. For example, given the question “The writer of the novel “1984” died in which continent? ”, MeLLo is unable to generate the helpful sub-question to find the final correct answer. In contrast, our DeepEdit, based on the constraints of [Conciseness], [Coherence], and [Receptiveness], correctly find the relevant new knowledge “Bucharest is located in the continent of North America.” and obtain the correct answer North America with respect to the new knowledge.
Taking a closer look, we observe that our DeepEdit’s induced reasoning is much more succinct than MeLLo, since we effectively filter out the irrelevant and redundant information during decoding. Because longer and noisy textual information can lead to more serious hallucinations and higher difficulty of reasoning, our DeepEdit exhibits a stronger capability of KE for LLMs. Overall, our DeepEdit method enforces LLMs’ outputs to follow our decoding constraints so as to soundly incorporate new knowledge in reasoning, which leads to the correct predictions in Figure 8.
6 Conclusion
We explore a new paradigm for knowledge editing of black-box LLMs: decoding with constraints. We have designed a new decoding method DeepEdit to improve the knowledge editing of LLMs. We enforce the LLM’s outputs to satisfy the desired properties with decoding at the reasoning step level. Based on the DeepEdit, we can improve the LLMs’ reasoning with new knowledge though higher conciseness, coherence, pertinence, and receptiveness. Our DeepEdit controls LLMs’ reasoning to soundly incorporate new knowledge by satisfying our constraints. In addition to DeepEdit, we provide two new benchmarks: MQuAKE-2002 and MQuAKE-hard to provide more precise and challenging evaluation for KE methods respectively. Extensive experiments demonstrate the effectiveness of our DeepEdit. Future work includes analyzing the effects of other constraints that can improve knowledge editing for large language models.